# SDFs from Unoriented Point Clouds using Neural Variational Heat Distances

### v2

@@ 1.1 Related Work
Figure 3: Comparison to grid-based method [15].
Using the heat equation for computing approximate geodesic distances has been introduced to geometry processing by Crane et al. [8, 9]. This method is based on two main steps. First, diffusing heat for a short time step starting from heat concentrated at the source points. Then, the normalized gradients of the heat function are used to compute the approximate geodesic distances by solving a Poisson equation. Many generalizations and improvements have been proposed, the latest being generalization to signed distance functions (GSD) [15] by diffusing signed normals. To the best of our knowledge, heat-based distance approximations have not been used in the neural setting so far. Following the original heat method, we also diffuse a scalar function from the input sources. However, we incorporate additional constraints in the second step, which allows us to correctly orient the heat gradients. Hence, unlike GSD, we are not diffusing normals but start with unoriented point clouds to compute a signed distance field. A comparison of our neural approach with $\approx 64^{3}$ parameters with the results of the GSD method [15] for two different grid sizes is provided in the inset Figure˜3. Note that for a grid size of $64$ (middle), where the number of variables corresponds to our number of parameters, the GSD reconstruction has considerable artifacts (e.g. fingers merging). To obtain a reconstruction comparable to ours, a larger grid size is required (right).
@@ 2.1 Unoriented Point Cloud to Unoriented Normals
which we assume to be well-defined almost everywhere, yields for smooth surfaces $\mathcal{S}$ a good approximation for the gradient of the distance function, i.e. the unoriented normal field of $\mathcal{S}$, which is well-defined up to the medial axis. However, note that this gradient is consistent only in the near field of the surface and corresponds to the unsigned distance and not the signed we actually would like to compute (Figure˜4).
@@ 2.2 Unoriented Normals to SDF
In the second step, we assume that we are given an approximate, unoriented normal field $n^{\tau}$ and aim for the SDF $\phi$ of the surface. To this end, we pick up the ansatz by Crane et al. [8] and consider a modification of the second step of their heat method. We construct an SDF approximation from $n^{\tau}$ with $\phi<0$ inside the surface, where the gradient should point in the opposite direction of $n^{\tau}$, and conversely, $\phi>0$ outside the surface with the gradient pointing in the same direction as $n^{\tau}$ (Figure˜5). To this end, we seek a minimizer of the normal fitting objective
@@
$\phi$
$u^{\tau}$
$\nabla\phi$
$n^{\tau}$
$n^{\tau}=-\frac{\nabla u^{\tau}}{\lvert\nabla u^{\tau}\rvert}$
$\nabla\phi\cdot n^{\tau}<0$
$\nabla\phi\cdot n^{\tau}>0$
@@ 3.1.1 Quadrature & Point Clouds
as a density rescaled approximation of the integral mean of $u$ on $\mathcal{S}$. In practice, $\epsilon$ is chosen such that every $\epsilon$-ball centered at any point contains at least 12 neighboring points. In Figure˜6, we show the heat solution $u^{\tau}$ for a non uniform input point cloud comparing uniform weights ˜11 with the local adaptive weights ˜12. It demonstrates that a proper weighting is necessary for non-uniform point clouds to obtain accurate normal directions from the heat solution.
@@ 3.1.2 Far Field Normals
The results in Figure˜7 indeed show that using the blended normal field leads to much better gradient norms and level sets farther from the surface. However, they degrade near the medial axis of the shapes. Thus creating a trade-off. If one is only interested in accurate results in a narrow band, the near field is sufficient. For our quantitative evaluations in the comparison section, we have always used the far field approach.
@@ 3.2.1 Inside & Outside
We construct the sets $\mathcal{B}^{-}$ on the inside of $\mathcal{S}$ and $\mathcal{B}^{+}$ on the outside using the following simple and efficient algorithm (see Figure˜8). We consider a regular grid with edge length $h>0$ and cell centers $\alpha h$ with $\alpha\in\mathbb{Z}^{3}$ and denote the corresponding open cells by $B_{\alpha}$. Step 1. We mark all cells whose closure has a non-empty intersection with $\mathcal{S}$ (i.e., containing points of the given point cloud) as interfacial cells. Step 2. Beginning with one cell intersecting the boundary of $\Omega$ that is not marked as interfacial, we successively mark each of the six neighboring cells for every outside cell $B_{\alpha}$, which are not yet marked, as outside. Step 3. We mark all remaining cells as interior. Finally, we define $\mathcal{B}^{+}$ as the union of all outside cells and $\mathcal{B}^{-}$ as the union of all inside cells. In practice, a rather coarse grid is sufficient, e.g., we used $64^{3}$ grid points and grid size $h=0.0375$. We note that for a very sparse point cloud, we may require a larger $h$. For example, for the sparse bean model in Table˜1, we used $2h$ as the grid size, which was the only case where different parameters were needed.
@@ 3.2.2 Smoothed Characteristic Functions
The normal approximation objective $\mathcal{E}_{n}$} depends on the sets $\phi<0]$ and $\phi>0]$, which means that it depends on the function $\phi$ in a non-smooth manner due to the thresholding operation. To facilitate the numerical minimization of $\mathcal{E}_{n}$}, we introduce a smooth approximation of the corresponding characteristic functions such that the objective becomes differentiable (see Figure˜9). To this end, we define $\eta_{\delta}(s)\coloneqq\eta(\delta^{-1}s)$ using the $C^{1}$ smooth, cubic Hermite blending function $\eta\colon\mathbb{R}\to\mathbb{R}_{\geq 0}$ with $\eta(s)=1$ for s<1s<-1, $\eta(s)=\tfrac{1}{4}(s+2)(s-1)^{2}$ for $|s|\leq 1$ and $\eta(s)=0$ for $s>1$. For $\delta\to 0$ and $s\neq 0$, $\eta_{\delta}(s)$ converges to $\chi_{\{s<0\}}$ and thus $\eta_{\delta}(\phi(\cdot))$ converges pointwise to $\chi_{[\phi<0]}$ on $\phi\neq 0]$ for a fixed $\phi$. Finally, this allows us to introduce the smoothed approximation
@@ 4 Experimental Results
We implemented our method using the PyTorch framework. All experiments took under an hour, including both optimization steps. The code is available at https://github.com/sweidemaier/HeatSDF
@@ 4.1 Parameters and Error Metrics
We have a single weighting parameter $\lambda_{\mathrm{fit}}$ in front of $\mathcal{E}_{\mathrm{fit}}$}, constraining the SDF to be close to zero on the input point cloud. We see a trade-off between a good surface fitting minimizing $\mathcal{E}_{\mathrm{fit}}$} and the normal alignment loss $\mathcal{E}_{n}$}. This is reflected in the decreasing reconstruction error for increasing $\lambda_{\mathrm{fit}}$ which comes with an increase of the eikonal error, see Figure˜10. Extremely high values of $\lambda_{\mathrm{fit}}$ might lead to surfaces with small artifacts. However, for the range covered in the plot, all surfaces can be recovered without artifacts. In all our experiments, we choose $\lambda_{\mathrm{fit}}=100$ as our focus lies on good SDF qualities.
@@
For all experiments, we choose $\tau=0.005$ for the small time step and $\hat{\tau}=0.1$. We use a smoothed blending function $\eta_{\delta}$ to approximate the characteristic functions of the sets $\phi<0]$ and $\phi>0]$ in (14). We study the dependence of our result on the parameter $\delta$, which corresponds to the steepness of $\eta_{\delta}$. For $\delta\in[0.001,0.01]$ the $L^{2}$ surface error only changes by orders of $10^{-6}$ and the eikonal error in a narrow band by $10^{-4}$, see Table˜2. Hence, the choice of this parameter is not very sensitive. In all our experiments, we choose $\delta=0.005$.
@@ 4.2 Comparisons
In Figure˜11, we show a quantitative comparison of the different methods over a range of shapes (cf. Appendix˜B, Figure˜17). We plot the surface reconstruction error $\mathbf{E_{recon}^{\mathcal{S}}}$ against the two measures for SDF quality $\mathbf{E_{eik}}$ and $\mathbf{E_{SDF}}$. The results show that our method achieves a balanced trade-off between an accurate surface reconstruction and a good SDF approximation. This is in contrast to HESS, which produces good reconstruction results but has larger eikonal and SDF errors and 1-Lip, which has low eikonal errors but loses detail in the surface reconstruction. Some results of SALD and HESS had flipped signs, which we inverted before measuring the errors.
Figure˜12 shows a qualitative comparison between the zero-level sets extracted with Marching Cubes for the different methods. Moreover, we plot for points, sampled in a narrow band near the surface, the neural distances compared the ground truth signed distance to the mesh. While HESS, HotSpot, and our method all achieve an accurate shape reconstruction, our method leads to the most accurate SDF estimation. We also note that HESS has a very accurate SDF estimation only very close to the surface and $1$-Lip consistently underestimates the distance, as expected from these methods. Those observations are also reflected in the corresponding table.
Our quadrature for the surface integral ˜12 accounting for the input point cloud density allows us to compute accurate SDFs even for spatially varying point densities. To demonstrate this, we consider a capped torus shape that can be described as the zero-level of a closed-form SDF [29]. We consider a dense, a sparse, and a non-uniform point cloud as well as an example with noise added to the surface; see Table˜1 for a visualization of the point clouds and quantitative values comparing the uniform and non-uniform case. All methods can deal with noisy and sparse input data. We provide the quantitative evaluation in the appendix Table˜3. However, HotSpot and SALD have difficulties dealing with the highly non-uniform point cloud, mixing inside and outside of the shape. In contrast, our SDF error only changes by a magnitude of $10^{-4}$. In Figure˜13, we show scatter plots for the SDF evaluation on the non-uniform capped torus.
Figure 11: Quantitative evaluation for different methods applied to various models (cf. Appendix˜B, Figure˜17), illustrating that our method effectively balances surface reconstruction and SDF approximation; top: eikonal error $\mathbf{E_{eik}}$ against surface error $\mathbf{E_{recon}^{\mathcal{S}}}$; bottom: SDF error $\mathbf{E_{SDF}}$ against surface error $\mathbf{E_{recon}^{\mathcal{S}}}$.
SALD $5.74\cdot 10^{-5}$ 0.1003 0.10630 0.01308
HESS $1.27\cdot 10^{-6}$ 0.6704 0.00134 0.02547
1-Lip $1.20\cdot 10^{-4}$ 0.0363 0.03793 0.01981
HotSpot $\mathbf{7.49\cdot 10^{-7}}$ 0.0651 0.00152 0.02578
Ours $1.59\cdot 10^{-6}$ 0.0901 0.00262 0.01184
@@
Figure 13: SDF error scatter plots for the non-uniformly sampled capped torus from Table˜1. Only 1-Lip and our method show a globally high consistency of the SDF. All other methods significantly deviate from the ground truth in the narrow band.
$\mathbf{\delta}$ $\mathbf{E_{recon}^{\mathcal{S}}}$ $\mathbf{E_{eik}}$
0.01 $1.312\cdot 10^{-5}$ 0.07020
0.0075 $1.287\cdot 10^{-5}$ 0.07133
0.005 $1.287\cdot 10^{-5}$ 0.07210
0.0025 $1.248\cdot 10^{-5}$ 0.07381
0.001 $1.161\cdot 10^{-5}$ 0.07418
@@ 4.3 Application: Neural PDE Heat Flow on Neural Surfaces
Let $\phi$ be the SDF of a surface $\mathcal{S}$ and consider the set of parallel surfaces
$\mathcal{S}_{c}\coloneqq\{x\in\Omega\mid\phi(x)=c\}$ for different values $c$ of the level set function, with $\mathcal{S}=\mathcal{S}_{0}$. Then $n(x)=\nabla\phi(x)$ is the normal at the point $x$ on $\mathcal{S}_{\phi(x)}$ and $P(x)=\text{Id}-n(x)\otimes n(x)$ the projection onto the tangent space $T_{x}\mathcal{S}_{\phi(x)}$. For the tangential gradient of a function $w$ one obtains $\nabla_{\mathcal{S}_{\phi(x)}}w(x)=(P\nabla w)(x)$.
@@
Altogether, a time-discrete, fully implicit Euler scheme can be formulated as a variational problem and thus solved numerically using a neural network approach. Figure˜14 (bottom) shows the results of neural heat diffusion on the neural zero-level set of the Max Planck head, in comparison to a piecewise affine, continuous finite element method on the original triangulation (top) from which the point cloud was extracted. Additionally, we present results of the finite element heat flow computation on the mesh extracted from the neural zero-level set (middle). We observe that the three approaches lead to very similar results qualitatively, indicating that our neural-PDE-on-neural-surface framework yields valid results without requiring to extract the mesh of the neural level set.

### v3

@@ 1.1 Related Work
Using the heat equation for computing approximate geodesic distances has been introduced to geometry processing by Crane et al. [8, 9]. This method is based on two main steps. First, diffusing heat for a short time step starting from heat concentrated at the source points. Then, the normalized gradients of the heat function are used to compute the approximate geodesic distances by solving a Poisson equation. Many generalizations and improvements have been proposed, the latest being generalization to signed distance functions (GSD) [15] by diffusing signed normals. To the best of our knowledge, heat-based distance approximations have not been used in the neural setting so far. Following the original heat method, we also diffuse a scalar function from the input sources. However, we incorporate additional constraints in the second step, which allows us to correctly orient the heat gradients. Hence, unlike GSD, we are not diffusing normals but start with unoriented point clouds to compute a signed distance field. A more detailed comparison to GSD can be found in Section˜4.2.2.
@@ 2.1 Unoriented Point Cloud to Unoriented Normals
which we assume to be well-defined almost everywhere, yields for smooth surfaces $\mathcal{S}$ a good approximation for the gradient of the distance function, i.e. the unoriented normal field of $\mathcal{S}$, which is well-defined up to the medial axis. However, note that this gradient is consistent only in the near field of the surface and corresponds to the unsigned distance and not the signed we actually would like to compute (Figure˜3).
@@ 2.2 Unoriented Normals to SDF
In the second step, we assume that we are given an approximate, unoriented normal field $n^{\tau}$ and aim for the SDF $\phi$ of the surface. To this end, we pick up the ansatz by Crane et al. [8] and consider a modification of the second step of their heat method. We construct an SDF approximation from $n^{\tau}$ with $\phi<0$ inside the surface, where the gradient should point in the opposite direction of $n^{\tau}$, and conversely, $\phi>0$ outside the surface with the gradient pointing in the same direction as $n^{\tau}$ (Figure˜4). To this end, we seek a minimizer of the normal fitting objective
@@ 3.1.1 Quadrature & Point Clouds
as a density rescaled approximation of the integral mean of $u$ on $\mathcal{S}$. In practice, $\epsilon$ is chosen such that every $\epsilon$-ball centered at any point contains at least 12 neighboring points. In Figure˜5, we show the heat solution $u^{\tau}$ for a non uniform input point cloud comparing uniform weights ˜11 with the local adaptive weights ˜12. It demonstrates that a proper weighting is necessary for non-uniform point clouds to obtain accurate normal directions from the heat solution.
@@ 3.1.2 Far Field Normals
The results in Figure˜6 indeed show that using the blended normal field leads to much better gradient norms and level sets farther from the surface. However, they degrade near the medial axis of the shapes. Thus creating a trade-off. If one is only interested in accurate results in a narrow band, the near field is sufficient. For our quantitative evaluations in the comparison section, we have always used the far field approach.
@@ 3.2.1 Inside & Outside
We construct the sets $\mathcal{B}^{-}$ on the inside of $\mathcal{S}$ and $\mathcal{B}^{+}$ on the outside using the following simple and efficient algorithm (see Figure˜7). We consider a regular grid with edge length $h>0$ and cell centers $\alpha h$ with $\alpha\in\mathbb{Z}^{3}$ and denote the corresponding open cells by $B_{\alpha}$. Step 1. We mark all cells whose closure has a non-empty intersection with $\mathcal{S}$ (i.e., containing points of the given point cloud) as interfacial cells. Step 2. Beginning with one cell intersecting the boundary of $\Omega$ that is not marked as interfacial, we successively mark each of the six neighboring cells for every outside cell $B_{\alpha}$, which are not yet marked, as outside. Step 3. We mark all remaining cells as interior. Finally, we define $\mathcal{B}^{+}$ as the union of all outside cells and $\mathcal{B}^{-}$ as the union of all inside cells. In practice, a rather coarse grid is sufficient, e.g., we used $64^{3}$ grid points and grid size $h=0.0375$. We note that for a very sparse point cloud, we may require a larger $h$. For example, for the sparse bean model in Table˜2, we used $2h$ as the grid size, which was the only case where different parameters were needed.
@@ 3.2.2 Smoothed Characteristic Functions
The normal approximation objective $\mathcal{E}_{n}$} depends on the sets $\phi<0]$ and $\phi>0]$, which means that it depends on the function $\phi$ in a non-smooth manner due to the thresholding operation. To facilitate the numerical minimization of $\mathcal{E}_{n}$}, we introduce a smooth approximation of the corresponding characteristic functions such that the objective becomes differentiable (see Figure˜8). To this end, we define $\eta_{\delta}(s)\coloneqq\eta(\delta^{-1}s)$ using the $C^{1}$ smooth, cubic Hermite blending function $\eta\colon\mathbb{R}\to\mathbb{R}_{\geq 0}$ with $\eta(s)=1$ for s<1s<-1, $\eta(s)=\tfrac{1}{4}(s+2)(s-1)^{2}$ for $|s|\leq 1$ and $\eta(s)=0$ for $s>1$. For $\delta\to 0$ and $s\neq 0$, $\eta_{\delta}(s)$ converges to $\chi_{\{s<0\}}$ and thus $\eta_{\delta}(\phi(\cdot))$ converges pointwise to $\chi_{[\phi<0]}$ on $\phi\neq 0]$ for a fixed $\phi$. Finally, this allows us to introduce the smoothed approximation
@@ 4 Experimental Results
We implemented our method using the PyTorch framework. All experiments took under an hour, including both optimization steps. All experiments were performed on a workstation with two AMD EPYC 7402 processors, 256GB RAM, and an NVIDIA A100 with 40GB memory. The code is available at https://github.com/sweidemaier/HeatSDF.
@@ 4.1 Parameters and Error Metrics
We have a single weighting parameter $\lambda_{\mathrm{fit}}$ in front of $\mathcal{E}_{\mathrm{fit}}$}, constraining the SDF to be close to zero on the input point cloud. We see a trade-off between a good surface fitting minimizing $\mathcal{E}_{\mathrm{fit}}$} and the normal alignment loss $\mathcal{E}_{n}$}. This is reflected in the decreasing reconstruction error for increasing $\lambda_{\mathrm{fit}}$ which comes with an increase of the eikonal error, see Figure˜9. Extremely high values of $\lambda_{\mathrm{fit}}$ might lead to surfaces with small artifacts. However, for the range covered in the plot, all surfaces can be recovered without artifacts. In all our experiments, we choose $\lambda_{\mathrm{fit}}=100$ as our focus lies on good SDF qualities.
@@
For all experiments, we choose $\tau=0.005$ for the small time step and $\hat{\tau}=0.1$. We use a smoothed blending function $\eta_{\delta}$ to approximate the characteristic functions of the sets $\phi<0]$ and $\phi>0]$ in (14). We study the dependence of our result on the parameter $\delta$, which corresponds to the steepness of $\eta_{\delta}$. For $\delta\in[0.001,0.01]$ the $L^{2}$ surface error only changes by orders of $10^{-6}$ and the eikonal error in a narrow band by $10^{-4}$, see Table˜1. Hence, the choice of this parameter is not very sensitive. In all our experiments, we choose $\delta=0.005$.
$\mathbf{\delta}$ $\mathbf{E_{recon}^{\mathcal{S}}}$ $\mathbf{E_{eik}}$
0.01 $1.312\cdot 10^{-5}$ 0.07020
0.0075 $1.287\cdot 10^{-5}$ 0.07133
0.005 $1.287\cdot 10^{-5}$ 0.07210
0.0025 $1.248\cdot 10^{-5}$ 0.07381
0.001 $1.161\cdot 10^{-5}$ 0.07418
@@ 4.2.1 Comparison to Neural SDF Methods
In Figure˜10, we show a quantitative comparison of the different methods over a range of shapes (cf. Appendix˜B, Figure˜18). We plot the surface reconstruction error $\mathbf{E_{recon}^{\mathcal{S}}}$ against the two measures for SDF quality $\mathbf{E_{eik}}$ and $\mathbf{E_{SDF}}$. The results show that our method achieves a balanced trade-off between an accurate surface reconstruction and a good SDF approximation. This is in contrast to HESS, which produces good reconstruction results but has larger eikonal and SDF errors and 1-Lip, which has low eikonal errors but loses detail in the surface reconstruction. Some results of SALD and HESS had flipped signs, which we inverted before measuring the errors.
Figure 10: Quantitative evaluation for different methods applied to various models (cf. Appendix˜B, Figure˜18), illustrating that our method effectively balances surface reconstruction and SDF approximation; top: eikonal error $\mathbf{E_{eik}}$ against surface error $\mathbf{E_{recon}^{\mathcal{S}}}$; bottom: SDF error $\mathbf{E_{SDF}}$ against surface error $\mathbf{E_{recon}^{\mathcal{S}}}$.
Figure˜11 shows a qualitative comparison between the zero-level sets extracted with Marching Cubes for the different methods. Moreover, we plot for points, sampled in a narrow band near the surface, the neural distances compared the ground truth signed distance to the mesh. While HESS, HotSpot, and our method all achieve an accurate shape reconstruction, our method leads to the most accurate SDF estimation. We also note that HESS has a very accurate SDF estimation only very close to the surface and $1$-Lip consistently underestimates the distance, as expected from these methods. Those observations are also reflected in the corresponding table.
Our quadrature for the surface integral ˜12 accounting for the input point cloud density allows us to compute accurate SDFs even for spatially varying point densities. To demonstrate this, we consider a capped torus shape that can be described as the zero-level of a closed-form SDF [29]. We consider a dense, a sparse, and a non-uniform point cloud as well as an example with noise added to the surface; see Table˜2 for a visualization of the point clouds and quantitative values comparing the uniform and non-uniform case. All methods can deal with noisy and sparse input data. We provide the quantitative evaluation in the appendix Table˜3. However, HotSpot and SALD have difficulties dealing with the highly non-uniform point cloud, mixing inside and outside of the shape. In contrast, our SDF error only changes by a magnitude of $10^{-4}$. In Figure˜12, we show scatter plots for the SDF evaluation on the non-uniform capped torus.
@@
SALD $5.74\cdot 10^{-5}$ 0.1003 0.10630 0.01308
HESS $1.27\cdot 10^{-6}$ 0.6704 0.00134 0.02547
1-Lip $1.20\cdot 10^{-4}$ 0.0363 0.03793 0.01981
HotSpot $\mathbf{7.49\cdot 10^{-7}}$ 0.0651 0.00152 0.02578
Ours $1.59\cdot 10^{-6}$ 0.0901 0.00262 0.01184
Figure 12: SDF error scatter plots for the non-uniformly sampled capped torus from Table˜2. Only 1-Lip and our method show a globally high consistency of the SDF. All other methods significantly deviate from the ground truth in the narrow band.
4.2.2 Comparison to Grid-based Methods
DoF $\mathbf{E_{recon}^{\mathcal{S}}}$ $\mathbf{E_{eik}}$ $\mathbf{E_{recon}^{n}}$ $\mathbf{E_{SDF}}$
GSD
$64^{3}$ $2.24\cdot 10^{-4}$ 0.0174 0.04005 0.01268
$128^{3}$ $5.95\cdot 10^{-5}$ 0.0087 0.01289 0.00661
$256^{3}$ $1.88\cdot 10^{-5}$ 0.0072 0.00497 0.00371
Ours $\sim 64^{3}$ $\mathbf{1.59\cdot 10^{-6}}$ 0.0901 0.00262 0.01184
Figure 13: Results for the same hand model as in Figure˜11. Our neural method outperforms the GSD approach [15] in terms of surface $L^{2}$ error ($E_{recon}^{\mathcal{S}}$}) and surface normal alignment ($E_{recon}^{n}$}) across all resolutions. Moreover, for a comparable number of degrees of freedom, our method achieves a similar accuracy with respect to the distance error ($E_{SDF}$}).
Figure 14: Comparison of our method with GSD on a lightbulb shape. Results are shown for GSD with approximately the same number of degrees of freedom (second from left) and for higher resolutions (two right columns), compared to our result (left).
In addition to the comparisons with state-of-the-art neural approaches presented in Section˜4.2.1, we also compare our method to the closest non-neural, heat-based approach Generalized Signed Distances (GSD) [15]. In GSD, the first step of the classical heat method is modified, such that the input geometry’s normal vectors are diffused using a short-time heat flow. The second step, namely normalizing those diffused vectors and integrating the resulting signed normal field to recover the signed distance function, remains unchanged. Unlike [15], we do not diffuse the normal field. Instead, we modify the second step of the heat method to obtain signed distances from an unsigned normal field. Nevertheless, both methods share the use of the heat flow to recover a continuous approximation of the normals of the distance function.
Figure˜13 shows a qualitative comparison between the zero-level sets for our neural method and that for GSD at three different grid resolutions, along with a quantitative evaluation. The grid-based GSD method with approximately the same number of degrees of freedom as our neural network exhibits notable artifacts in regions of near self-contact (e.g., merged fingers). Additionally, it struggles in regions of high curvature (e.g., the thin wires of the lightbulb shape), as illustrated in Figure˜14. Achieving a reconstruction quality comparable to ours requires a significantly higher grid resolution, i.e. more degrees of freedom. This is also reflected by the quantitative results in Figure˜10. Our method requires on average approximately $25$ minutes to train, whereas the grid-based method takes, on average across the presented models, $1.5$ minutes ($64^{3}$), $10$ minutes ($128^{3}$), and $75$ minutes ($256^{3}$) to compute a solution. For these comparisons, we used the open source code provided by the authors of [15] with some straightforward parallelizations. In both cases, runtimes heavily depend on the size of the input point cloud.
@@ 4.3 Application: Neural PDE Heat Flow on Neural Surfaces
Let $\phi$ be the SDF of a surface $\mathcal{S}$ and consider the set of parallel surfaces $\mathcal{S}_{c}\coloneqq\{x\in\Omega\mid\phi(x)=c\}$ for different values $c$ of the level set function, with $\mathcal{S}=\mathcal{S}_{0}$. Then $n(x)=\nabla\phi(x)$ is the normal at the point $x$ on $\mathcal{S}_{\phi(x)}$ and $P(x)=\text{Id}-n(x)\otimes n(x)$ the projection onto the tangent space $T_{x}\mathcal{S}_{\phi(x)}$. For the tangential gradient of a function $w$ one obtains $\nabla_{\mathcal{S}_{\phi(x)}}w(x)=(P\nabla w)(x)$.
@@
Altogether, a time-discrete, fully implicit Euler scheme can be formulated as a variational problem and thus solved numerically using a neural network approach. Figure˜15 (bottom) shows the results of neural heat diffusion on the neural zero-level set of the Max Planck head, in comparison to a piecewise affine, continuous finite element method on the original triangulation (top) from which the point cloud was extracted. Additionally, we present results of the finite element heat flow computation on the mesh extracted from the neural zero-level set (middle). We observe that the three approaches lead to very similar results qualitatively, indicating that our neural-PDE-on-neural-surface framework yields valid results without requiring to extract the mesh of the neural level set.